# Toward a test of Gaussianity of a gravitational wave background

### v1

@@ Appendix A Central Limit Theorem and Large Scale Modes
The central limit theorem and law of large numbers play an important role in cosmology Verde (2010); Trotta (2017). A sample (spatial) average of a summary statistic, constructed from a cosmological data set, is assumed to converge to the true mean of the underlying probability distribution of that statistic. Beyond that, we also often assume that the covariance associated with said measurement is Gaussian. For example, if we measure some summary statistic $u_{\mu}$ [^2] This could be the two-point correlation function, the Minkowski Functionals etc. We include a $\mu$ subscript to denote a possible binning of the statistic in density, separation etc. In this appendix we simply measure the cumulants of a field, which are scalars. then the sample average is taken to be an unbiased estimator of the ensemble average $\langle u_{\mu}\rangle$, and the uncertainty of the measurement is inferred from a covariance matrix $\Sigma_{\mu\nu}$. This can be constructed either by numerical estimation from mock data or explicitly calculating the ensemble average $\Sigma_{\mu\nu}\sim\langle u_{\mu}u_{\nu}\rangle$. By using the ensemble average $\langle u_{\mu}\rangle$ and covariance $\Sigma_{\mu\nu}$ to quantify the statistical properties of $u_{\mu}$, we are effectively modelling it as a multivariate Gaussian. In this appendix, we will consider the extent to which these approximations can be made.
We start with an all-sky map of a Gaussian random field $\Delta_{i}$ drawn from an angular power spectrum $C_{\ell}$, where $i$ denotes the pixel number on the two-sphere [^3] To generate and manipulate the maps we utilise Healpix (http://healpix.sourceforge.net) (Gorski et al., 2005).. We measure the cumulants of this field, because they constitute the simplest set of non-trivial summary statistics that can be used to present our point. We define
$\kappa_{n}\equiv{1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\Delta_{i}^{n}.$ (11)
@@
$P_{\Delta_{i}^{n}}(x)={|x|^{(1-n)/n}\over\sqrt{2\pi\sigma^{2}}}e^{-|x|^{2/n}/2\sigma^{2}},$ (12)
@@
$\kappa_{2}={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\Delta_{i}^{2}={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}y_{i}$ (13)
where $y_{i}$ are independent, $\chi^{2}(1)$ random variables and $\lambda_{i}$ are the eigenvalues of the pixel covariance matrix $\Sigma$. We have defined $\chi^{2}(n)$ as a chi square distribution with $n$ degrees of freedom. The probability distribution function (PDF) of $\kappa_{2}$ does not have a closed form solution for the general case in which $\lambda_{i}\neq\lambda_{j}$ for $i\neq j$, but it is straightforward to extract its moments from the second sum in equation (13) since $y_{i}$ are independent, identically distributed (iid). For example the mean and variance of $\kappa_{2}$ are given by
${\rm E}(\kappa_{2})$ $=$ ${1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}{\rm E}(y_{i})={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}$ (14)
${\rm var}(\kappa_{2})$ $=$ $\sum_{i=1}^{N_{\rm pix}}{\rm var}\left({\lambda_{i}\over N_{\rm pix}}y_{i}\right)={2\over N_{\rm pix}^{2}}\sum_{i=1}^{N_{\rm pix}}\lambda^{2}_{i}\,.$ (15)
@@
$\kappa_{2}={\sigma^{2}\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}y_{i}={\sigma^{2}\over N_{\rm pix}}\chi^{2}(N_{\rm pix})$ (16)
@@
$\Sigma_{ij}=\sum_{\ell}{2\ell+1\over 4\pi}C_{\ell}W_{\ell}^{2}(\theta_{G})P_{\ell}(x_{ij})$ (17)
where we smooth the field with a Gaussian kernel $W_{\ell}=\exp[-\ell(\ell+1)\theta_{G}^{2}/2]$ and angular smoothing scale $\theta_{G}=2\sqrt{4\pi/N_{\rm pix}}$. $P_{\ell}(x)$ are the Legendre polynomials and $x_{ij}=\hat{n}_{i}.\hat{n}_{j}$, where $\hat{n}_{i}$ is the unit vector pointing to pixel $i$ [^4] In equation (17) have used the fact that the data is an all-sky map, for point distributions or masked data a different estimator for $\Sigma_{ij}$ should be used.. The diagonal elements of $\Sigma_{ij}$ are
$\sigma^{2}\equiv\Sigma_{ii}=\sum_{\ell}{2\ell+1\over 4\pi}C_{\ell}W_{\ell}^{2}(\theta_{G})\,.$ (18)
@@
We observe that as the power spectrum $C_{\ell}$ becomes increasingly blue tilted, the eigenvalues become sharply peaked at $i\sim 1$. Because the eigenvalues act as coefficients multiplying the random variables $y_{i}$ in the expression (13) for $\kappa_{2}$, the presence of a hierarchy implies that the sum will be dominated by a small number of random variables. In this case, the large $N_{\rm pix}$ limit will not Gaussianize the $\kappa_{2}$ PDF. Indeed, as the hierarchy in $\lambda_{i}$ becomes increasingly pronounced, $\kappa_{2}$ will be approximated as being drawn from a $\chi^{2}(n)$ distribution with $n\sim{\cal O}(1)$. In Figure 7 we present the PDF of $\kappa_{2}$ for $\alpha=-4,-2,0,2$. The blue histograms were obtained by generating $N_{\rm real}=10^{4}$ realisations of a Gaussian random field on the sphere for each $C_{\ell}=\ell^{\alpha}$ and estimating $\kappa_{2}$ by taking the pixel sum (13). The black dashed lines in Figure 7 are Gaussian distributions with mean and variance given by equations (14,15) and the vertical green solid/dashed lines are the mean and ±$\pm 2\sigma$ bounds of this Gaussian approximation. The blue vertical lines are the median and $95\%$ confidence region of the numerically reconstructed PDF.
@@
The higher order cumulants are less tractable than the variance, but the same logic applies. The skewness $\kappa_{3}$ involves triple products of the Gaussian random variable $\Delta_{i}$, which are less straightforward to de-correlate than the variance. However, the same issues arise – in the presence of large scale correlations in the data the probability distributions of the summary statistic is highly non-Gaussian. In Figure 8 we present the PDFs of $\kappa_{3}$ for $\alpha=-4,-2,0,+2$. As before, the solid/dashed blue lines are the median and $95\%$ confidence region of the numerically reconstructed PDF (cf blue histograms). The black dashed curve is a Gaussian distribution inferred from the ensemble average of the statistics Srednicki (1993); $\langle\kappa_{3}\rangle=0$ and
@@
Figure 8: Same as Figure 7, but for the cubic cumulant $\kappa_{3}$.

### v2

@@ Appendix A Central Limit Theorem and Large Scale Modes
The central limit theorem and law of large numbers play an important role in cosmology [106, 107]. A sample (spatial) average of a summary statistic, constructed from a cosmological data set, is assumed to converge to the true mean of the underlying probability distribution of that statistic. Beyond that, we also often assume that the covariance associated with said measurement is Gaussian. For example, if we measure some summary statistic $u_{\mu}$ [^3] This could be the two-point correlation function, the Minkowski Functionals etc. We include a $\mu$ subscript to denote a possible binning of the statistic in density, separation etc. In this appendix we simply measure the cumulants of a field, which are scalars. then the sample average is taken to be an unbiased estimator of the ensemble average $\langle u_{\mu}\rangle$, and the uncertainty of the measurement is inferred from a covariance matrix $\Sigma_{\mu\nu}$. This can be constructed either by numerical estimation from mock data or explicitly calculating the ensemble average $\Sigma_{\mu\nu}\sim\langle u_{\mu}u_{\nu}\rangle$. By using the ensemble average $\langle u_{\mu}\rangle$ and covariance $\Sigma_{\mu\nu}$ to quantify the statistical properties of $u_{\mu}$, we are effectively modelling it as a multivariate Gaussian. In this appendix, we will consider the extent to which these approximations can be made.
We start with an all-sky map of a Gaussian random field $\Delta_{i}$ drawn from an angular power spectrum $C_{\ell}$, where $i$ denotes the pixel number on the two-sphere [^4] To generate and manipulate the maps we utilise Healpix (http://healpix.sourceforge.net) [108].. We measure the cumulants of this field, because they constitute the simplest set of non-trivial summary statistics that can be used to present our point. We define
$\kappa_{n}\equiv{1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\Delta_{i}^{n}.$ (A.1)
@@
$P_{\Delta_{i}^{n}}(x)={|x|^{(1-n)/n}\over\sqrt{2\pi\sigma^{2}}}e^{-|x|^{2/n}/2\sigma^{2}},$ (A.2)
@@
$\kappa_{2}={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\Delta_{i}^{2}={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}y_{i}$ (A.3)
where $y_{i}$ are independent, $\chi^{2}(1)$ random variables and $\lambda_{i}$ are the eigenvalues of the pixel covariance matrix $\Sigma$. We have defined $\chi^{2}(n)$ as a chi square distribution with $n$ degrees of freedom. The probability distribution function (PDF) of $\kappa_{2}$ does not have a closed form solution for the general case in which $\lambda_{i}\neq\lambda_{j}$ for $i\neq j$, but it is straightforward to extract its moments from the second sum in equation (A.3) since $y_{i}$ are independent, identically distributed (iid). For example the mean and variance of $\kappa_{2}$ are given by
${\rm E}(\kappa_{2})$ $=$ ${1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}{\rm E}(y_{i})={1\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}\lambda_{i}$ (A.4)
${\rm var}(\kappa_{2})$ $=$ $\sum_{i=1}^{N_{\rm pix}}{\rm var}\left({\lambda_{i}\over N_{\rm pix}}y_{i}\right)={2\over N_{\rm pix}^{2}}\sum_{i=1}^{N_{\rm pix}}\lambda^{2}_{i}\,.$ (A.5)
@@
$\kappa_{2}={\sigma^{2}\over N_{\rm pix}}\sum_{i=1}^{N_{\rm pix}}y_{i}={\sigma^{2}\over N_{\rm pix}}\chi^{2}(N_{\rm pix})$ (A.6)
@@
$\Sigma_{ij}=\sum_{\ell}{2\ell+1\over 4\pi}C_{\ell}W_{\ell}^{2}(\theta_{G})P_{\ell}(x_{ij})$ (A.7)
where we smooth the field with a Gaussian kernel $W_{\ell}=\exp[-\ell(\ell+1)\theta_{G}^{2}/2]$ and angular smoothing scale $\theta_{G}=2\sqrt{4\pi/N_{\rm pix}}$. $P_{\ell}(x)$ are the Legendre polynomials and $x_{ij}=\hat{n}_{i}.\hat{n}_{j}$, where $\hat{n}_{i}$ is the unit vector pointing to pixel $i$ [^5] In equation (A.7) have used the fact that the data is an all-sky map, for point distributions or masked data a different estimator for $\Sigma_{ij}$ should be used.. The diagonal elements of $\Sigma_{ij}$ are
$\sigma^{2}\equiv\Sigma_{ii}=\sum_{\ell}{2\ell+1\over 4\pi}C_{\ell}W_{\ell}^{2}(\theta_{G})\,.$ (A.8)
@@
We observe that as the power spectrum $C_{\ell}$ becomes increasingly blue tilted, the eigenvalues become sharply peaked at $i\sim 1$. Because the eigenvalues act as coefficients multiplying the random variables $y_{i}$ in the expression (A.3) for $\kappa_{2}$, the presence of a hierarchy implies that the sum will be dominated by a small number of random variables. In this case, the large $N_{\rm pix}$ limit will not Gaussianize the $\kappa_{2}$ PDF. Indeed, as the hierarchy in $\lambda_{i}$ becomes increasingly pronounced, $\kappa_{2}$ will be approximated as being drawn from a $\chi^{2}(n)$ distribution with $n\sim{\cal O}(1)$. In Figure 7 we present the PDF of $\kappa_{2}$ for $\alpha=-4,-2,0,2$. The blue histograms were obtained by generating $N_{\rm real}=10^{4}$ realisations of a Gaussian random field on the sphere for each $C_{\ell}=\ell^{\alpha}$ and estimating $\kappa_{2}$ by taking the pixel sum (A.3). The black dashed lines in Figure 7 are Gaussian distributions with mean and variance given by equations (A.4,A.5) and the vertical green solid/dashed lines are the mean and ±$\pm 2\sigma$ bounds of this Gaussian approximation. The blue vertical lines are the median and $95\%$ confidence region of the numerically reconstructed PDF.
@@
Figure 8: Same as Figure 7, but for the cubic cumulant $\kappa_{3}$.
The higher order cumulants are less tractable than the variance, but the same logic applies. The skewness $\kappa_{3}$ involves triple products of the Gaussian random variable $\Delta_{i}$, which are less straightforward to de-correlate than the variance. However, the same issues arise – in the presence of large scale correlations in the data the probability distributions of the summary statistic is highly non-Gaussian. In Figure 8 we present the PDFs of $\kappa_{3}$ for $\alpha=-4,-2,0,+2$. As before, the solid/dashed blue lines are the median and $95\%$ confidence region of the numerically reconstructed PDF (cf blue histograms). The black dashed curve is a Gaussian distribution inferred from the ensemble average of the statistics [66]; $\langle\kappa_{3}\rangle=0$ and